The length of the top of the table is 7m greater than the width. The area is 78m^2. What is the dimensions of the table?
Answer by mananth(16949)

The length of the top of the table is 7m greater than the width. The area is 78m^2. What is the dimensions of the table?
...
let width be x m
length = x+7 m
..
L*W = Area
x(x+7)=78
x^2+7x=78
-78 -78
x^2+7x-78=0
x^2+13x-6x-78=0
x(x+13)-6(x+13)=0
(x+13)(x-6)=0
ignore negative
x= 6 m the width.
length = 6+7 = 13 m.
...
m.ananth@hotmail.ca